# Spectral Small-Incremental-Entangling: Breaking Quasi-Polynomial Complexity Barriers in Long-Range Interacting Systems

### v1

@@ VI.4 Proof of Theorem 3
$=\sum_{s=0}^{\infty}z^{s}\int_{0}^{1}dx_{1}\int_{0}^{x_{1}}dx_{2}\cdots\int_{0}^{x_{s-1}}dx_{s}V_{AB}(x_{1})V_{AB}(x_{2})\cdots V_{AB}(x_{s}).$ (169)
where $V_{AB}(x)=e^{-xzH_{0}}V_{AB}e^{xzH_{0}}$ and $z\in\mathbb{C}$. First, we prove that for small $|z|$ the operator $\Psi$ can be approximated by another operator $\tilde{\Psi}_{\epsilon}$ with a small Schmidt rank:
@@
$\left\|\Psi-\tilde{\Psi}_{\epsilon}\right\|\leq\epsilon$ (171)
@@
We postpone the proof of the proposition to Sec. VI.4.1 below. Using this proposition, we now give a proof of the main statement, presented in the following.
We begin with the real-time evolution $t=mt_{0}$ with $t_{0}$ satisfying the condition (170). We generally obtain
@@
$\left\|e^{-imt_{0}(H_{0}+V_{AB})}-\left(e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right)^{m}\right\|\leq\sum_{j=0}^{m-1}\left\|e^{-it_{0}(H_{0}+V_{AB})}-e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|\cdot\left\|e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|^{j},$ (174)
@@
$\left\|e^{-it_{0}(H_{0}+V_{AB})}-e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|\leq\left\|\Psi-\tilde{\Psi}_{\epsilon}\right\|\leq\epsilon.$ (175)
@@
Next, we consider the imaginary-time evolution $\beta=2m\beta_{0}$ with $\beta_{0}$ satisfying the condition (170). Using $\tilde{\Psi}_{\epsilon}$, we examine $e^{\beta_{0}H_{0}}\tilde{\Psi}_{\epsilon}$, which approximates $e^{\beta_{0}(H_{0}+V_{AB})}$ by
@@
In total, we obtain the following error bound:
$\left\|e^{\beta_{0}(H_{0}+V_{AB})}-e^{\beta_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|_{p}\leq e^{2\beta_{0}\tilde{g}}\epsilon\left\|e^{\beta_{0}(H_{0}+V_{AB})}\right\|_{p},$ (184)
@@
This proves the second main inequality in (3).
@@ VI.4.1 Proof of Proposition 12
We here introduce the decomposition of $V_{AB}$ as follows:
$V_{AB}=\sum_{m=0}^{\infty}V_{r_{m}}$ (188)
with $V_{r_{m}}$ defined as
$V_{r_{m}}:=\sum_{j\in[r_{m},r_{m+1})}V_{j},\quad r_{m}=4^{m/\kappa}$ (189)
Here, we adopt the notation $\overline{V_{r_{m}}}$ as
@@
For the second summation in (VI.4.1), we can derive
@@
$\leq\left(\frac{8C_{0}\tilde{g}}{3}\right)^{s}\sum_{\bar{m}=M+1}^{\infty}2^{-\bar{m}-1}\leq 2^{-M-1}(3C_{0}\tilde{g})^{s}$ (201)
@@
Finally, for the third summation in (VI.4.1), we can derive
@@
${\rm SR}\left(\Psi_{Q,M,s_{0}}\right)\leq\sum_{s=0}^{s_{0}}\sum_{m_{1}+m_{2}+\cdots+m_{s}\leq M}\sum_{q_{1}+q_{2}+\cdots+q_{s}\leq Q}\prod_{j=1}^{s}{\rm SR}\left[{\rm ad}_{H_{0}}^{q_{j}}(V_{r_{m_{j}}})\right]$ (206)
@@
By applying the inequality (210) to (206), we obtain
@@
Note that the last equation comes from the Hockey-stick identity.
@@ VII Generalized entanglement area law
The motivation in this section is to figure out the rigorous condition such that the generalized area law is recovered (Fig. 3). Here, we impose the following additional assumptions, which are typically considered for the area-law proof under the adiabatic path [53, 26, 22, 54]. We here restrict ourselves to the following boundary-adiabatic path:
@@
Consider a family of Hamiltonians parametrized by $\nu$ through the boundary interaction $V_{AB}^{(\nu)}$:
@@
Moreover, in a more general setting, one can consider an arbitrary parameterization of the Hamiltonian like $H_{A}^{(\nu)}+V_{AB}^{(\nu)}+H_{B}^{(\nu)}$. However, it is likely that we may find a bad adiabatic path to connect the initial ground state to a highly entangled ground state as shown in Ref. [49]. Although we have not found an explicit counterexample, we conjecture that only the standard adiabatic path condition alone is insufficient to ensure the generalized area law.
@@
Let us denote the initial-state entanglement by using the $\infty$-Rényi entanglement:
@@
Also, the entanglement entropy $E_{\alpha=1}(\Omega^{(1)})$ for the ground state $|\Omega^{(1)}\rangle$ is upper-bounded by
@@
In our setup, the boundary size is roughly estimated by $\tilde{g}/\Delta$ since $\tilde{g}$ characterizes the dynamical entanglement rate from the spectral SIE. Indeed, if we consider a graph system with finite-range interactions, $\tilde{g}$ is indeed proportional to the boundary size between $A$ and $B$. From this perspective, our result qualitatively yields
@@ VII.1 Approximate-Ground-Space-Projection (AGSP)
In the proof of Theorem 4, we combine the quantum adiabatic theorem [55] with the so-called approximate ground-state projection (AGSP) approach [39, 40, 44, 43]
@@
    In the standard AGSP framework, the primary focus is on bounding the Schmidt rank of the AGSP operator.
@@
There exists an AGSP operator $K_{\beta}$ with the following properties:
@@
where $|\phi\rangle$ is an arbitrary product state between $A$ and $B$, and $|\phi_{D}\rangle$ is an appropriate quantum state with a Schmidt rank up to $D$.
@@
Here, the AGSP $K_{\beta}$ itself may not be approximated by an operator with the Schmidt rank $D$ (see also Sec. VI). However, in applying the AGSP formalism to the ground state, the Schmidt rank approximation with respect to $K_{\beta}|\phi\rangle$ is more crucial. We utilize this proposition to achieve the ground state approximation in Theorem 4, as well as in Theorem 5 below, concerning the polynomial complexity of the ground state in 1D long-range interacting systems.
@@ VII.1.1 Proof of Proposition 14
For the proof, we employ the AGSP operator $K_{\beta}$ in the form
@@
Next, we estimate the SE strength $\bar{{\mathcal{J}}}(K_{\beta})$ based on Definition 1 to prove the inequality (230). We use Lemma 4 and obtain the inequality of
@@ VII.2 Proof of Theorem 4
We begin with considering the Schmidt-rank truncation problem for the initial ground state $|\Omega^{(\nu)}\rangle$ with $\nu=0$. In general, any restrictions to Rényi entanglement with $\alpha\geq 1$ cannot ensure the efficiency guarantee of the truncation [30, 31]. However, by utilizing the fact that $|\Omega^{(0)}\rangle$ is the gapped ground state, we can prove the following proposition, which plays a key role in the proof:
@@
$|\Omega^{(\nu)}\rangle=\sum_{s}\lambda_{s}|\phi_{A,s}\rangle\otimes|\phi_{B,s}\rangle$ (241)
We then derive the following general equation:
@@
where we use $\log\left(\lambda^{2}_{1}\right)=E_{\infty}(\Omega^{(\nu)})=S_{0}$. Only from the information on the $\infty$-Rényi entanglement, we cannot get further information on $\lambda_{s}$ with $s\geq 2$.
To overcome the difficulty, we recover the decay rate of $\{\lambda_{s}\}_{s}$ from the condition (242) using the AGSP formalism. We now decompose $|\phi_{A,1},\phi_{B,1}\rangle$ in Eq. (241) into
@@
where $\langle\Omega^{(\nu)}|\Omega_{\bot}^{(\nu)}\rangle=0$. To recover $|\Omega^{(\nu)}\rangle$ from $|\phi_{A,1},\phi_{B,1}\rangle$, we utilize the AGSP $K_{\beta}$ constructed in Proposition 14, which gives
@@
where we use the parameter $\tilde{g}$ that upper-bounds the SE strength as in (218). Therefore, by letting $|\psi_{D}\rangle:=\lambda_{1}^{-1}|\phi_{D}\rangle$ and combining the inequalities (VII.2) and (245), we obtain
$\left\||\Omega^{(\nu)}\rangle-|\psi_{D}\rangle\right\|$ $\leq\frac{3}{\lambda_{1}}e^{-\beta\Delta^{2}}+\frac{e^{2\beta\Delta\tilde{g}}}{\lambda_{1}D^{1/2}}.$ (246)
By choosing $\beta$ such that $e^{-\beta\Delta^{2}}=e^{2\beta\Delta\tilde{g}}D^{-1/2}$, or
@@
where we use Eq. (242) and the notation $\kappa_{\Delta}$ in Eq. (240). This completes the proof. $\square$
@@
To analyze the target ground state $|\Omega^{(\nu)}\rangle$ with $\nu=1$., we here utilize the adiabatic time evolution using the parametrized Hamiltonian $H^{(\nu)}$ ($0\leq\nu\leq 1$) as follows:
@@
From Ref. [55], we have the approximation error as
@@
We next estimate the lower bound of $a_{\ell}$. From the Eckart–Young theorem [72], by employing the quantum state $|\psi_{D}\rangle$ in Proposition 15, we obtain
@@
By combining the error bound (251) and Eq. (254) with $a_{\ell}^{2}\geq 1/2$, we derive the inequality of
@@
We then construct the approximation of the target ground state $|\Omega^{(1)}\rangle$ based on the state $|\psi_{\ell,1/\epsilon}\rangle$. For this purpose, we first decompose
@@
$\sum_{s=1}^{\infty}\tilde{\lambda}_{\beta,s}\leq\sum_{s=1}^{\infty}\tilde{\lambda}_{s}\bar{{\mathcal{J}}}\left(K_{\beta}/c_{1}\right)$ $=\bar{{\mathcal{J}}}\left(K_{\beta}/c_{1}\right)e^{E_{\alpha=1/2}(\psi_{1,1/\varepsilon})/2}\leq 3\ell^{1/2}e^{2\beta\Delta\tilde{g}+\tilde{g}/\varepsilon},$ (266)
@@
$\left\|c_{1}^{-1}K_{\beta}|\psi_{\ell,t}\rangle-|\psi_{\beta,D}\rangle\right\|\leq\frac{3\ell^{1/2}e^{(2\beta\Delta+1/\varepsilon)\tilde{g}}}{D^{1/2}},$ (267)
where we use a similar inequality to (33).
@@
By choosing the parameter $\beta$ such that $e^{-\beta\Delta^{2}}=e^{2\beta\Delta\tilde{g}}D^{-1/2}$, or
@@
Finally, we upper-bound the coefficient $12\ell^{1/2}e^{\tilde{g}/\varepsilon}$ in Eq. (270). We have set in Eqs. (257) and (260)
@@
$\left\||\psi\rangle-|\psi_{D}\rangle\right\|\leq\frac{C}{D^{\kappa}}\quad\textrm{for}\quad\forall D\in\mathbb{N},$ (273)
where $C$ and $\kappa$ are positive constant. Then, the entanglement entropy is upper-bounded by
@@
$c_{\kappa,1}:=\frac{2-2^{-2\kappa}}{\kappa(1-2^{-2\kappa})},\quad c_{\kappa,2}:=\frac{(6+2\kappa)\log(2)}{(1-2^{-2\kappa})^{2}}$ (275)
By using this lemma with the first main inequality (222), i.e., $\kappa\to\kappa_{\Delta}$ and $C\to C_{\tilde{g},\Delta,S_{0}}$, we immediately prove the second main inequality (225). This completes the proof of Theorem 4. $\square$
Proof of Lemma 16. We here consider a set of $\{D_{p}\}_{p=1}^{\infty}$ such that $D_{p}=2^{p+p_{0}}$ and define $\delta_{p}$ as
@@
where $p_{0}$ is chosen appropriately afterward. Then, by using Ref. [44, Supplemental Proposition 3], we obtain
@@
$\leq p_{0}\log(2)+\frac{2^{-2p_{0}\kappa}C^{2}}{1-2^{-2\kappa}}\log\left(\frac{6\cdot 2^{p_{0}(2\kappa+1)}}{C^{2}}\right)+\frac{C^{2}2^{-2\kappa(p_{0}+1)}(2\kappa+1)\log(2)}{(1-2^{-2\kappa})^{2}}$ (277)
@@
VIII Polynomial-time simulation of 1D long-range interacting systems

### v2

@@ VI.4 Proof of Theorem 3
$=\sum_{s=0}^{\infty}z^{s}\int_{0}^{1}dx_{1}\int_{0}^{x_{1}}dx_{2}\cdots\int_{0}^{x_{s-1}}dx_{s}V_{AB}(x_{1})V_{AB}(x_{2})\cdots V_{AB}(x_{s}),$ (169)
where $V_{AB}(x)=e^{-xzH_{0}}V_{AB}e^{xzH_{0}}$ and $z\in\mathbb{C}$. Using it, we have
$e^{z(H_{0}+V_{AB})}=e^{zH_{0}}\Psi.$ (170)
We notice that $e^{zH_{0}}$ is a product operator between $A$ and $B$. First, we prove that for small $|z|$ the operator $\Psi$ can be approximated by another operator $\tilde{\Psi}_{\epsilon}$ with a small Schmidt rank:
@@
$\left\|\Psi-\tilde{\Psi}_{\epsilon}\right\|\leq\epsilon,$ (172)
@@
We defer the proof of the proposition to Sec. VI.4.1 below. Using this proposition, we now give a proof of the main statement, as follows.
We first consider the real-time evolution z=itz=-it ($t\in\mathbb{R}$) and decompose the total time $t$ into $mt_{0}$ with $t_{0}$ satisfying the condition (171). We generally obtain
@@
$\left\|e^{-imt_{0}(H_{0}+V_{AB})}-\left(e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right)^{m}\right\|\leq\sum_{j=0}^{m-1}\left\|e^{-it_{0}(H_{0}+V_{AB})}-e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|\cdot\left\|e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|^{j}.$ (175)
@@
$\left\|e^{-it_{0}(H_{0}+V_{AB})}-e^{-it_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|\leq\left\|\Psi-\tilde{\Psi}_{\epsilon}\right\|\leq\epsilon,$ (176)
@@
Next, we consider the imaginary-time evolution $z=\beta$ with the decomposition $z=2m\beta_{0}$, where $\beta_{0}$ satisfies the condition (171). Using $\tilde{\Psi}_{\epsilon}$, we examine $e^{\beta_{0}H_{0}}\tilde{\Psi}_{\epsilon}$, which approximates $e^{\beta_{0}(H_{0}+V_{AB})}$ by
@@
In total, we derive the following error bound:
$\left\|e^{\beta_{0}(H_{0}+V_{AB})}-e^{\beta_{0}H_{0}}\tilde{\Psi}_{\epsilon}\right\|_{p}\leq e^{2\beta_{0}\tilde{g}}\epsilon\left\|e^{\beta_{0}(H_{0}+V_{AB})}\right\|_{p}.$ (185)
@@
We thus prove the second main inequality in (3).
@@ VI.4.1 Proof of Proposition 12
We introduce the decomposition of $V_{AB}$ as follows:
$V_{AB}=\sum_{m=0}^{\infty}V_{r_{m}},$ (189)
with $V_{r_{m}}$ defined by
$V_{r_{m}}:=\sum_{j\in[r_{m},r_{m+1})}V_{j},\quad r_{m}=4^{m/\kappa}.$ (190)
Here, we introudce the notation $\overline{V_{r_{m}}}$ as
@@
For the second summation in (VI.4.1), we obtain
@@
$\leq\left(\frac{8C_{0}\tilde{g}}{3}\right)^{s}\sum_{\bar{m}=M+1}^{\infty}2^{-\bar{m}-1}\leq 2^{-M-1}(3C_{0}\tilde{g})^{s},$ (202)
@@
Finally, for the third summation in (VI.4.1), we deduce
@@
${\rm SR}\left(\Psi_{Q,M,s_{0}}\right)\leq\sum_{s=0}^{s_{0}}\sum_{m_{1}+m_{2}+\cdots+m_{s}\leq M}\sum_{q_{1}+q_{2}+\cdots+q_{s}\leq Q}\prod_{j=1}^{s}{\rm SR}\left[{\rm ad}_{H_{0}}^{q_{j}}(V_{r_{m_{j}}})\right].$ (207)
@@
By applying the inequality (211) to (207), we obtain the upper bound on the Schmidt rank as follows:
@@
Note that the last equation follows from the hockey-stick identity.
@@ VII Generalized entanglement area law
The motivation in this section is to figure out the rigorous condition such that the generalized area law is recovered (Fig. 3). Here, we impose the following additional assumptions, which are typically considered for the area-law proof under the adiabatic path [53, 26, 22, 54]. We restrict ourselves to the following boundary-adiabatic path:
@@
We consider a family of Hamiltonians parametrized by $\nu$ through the boundary interaction $V_{AB}^{(\nu)}$:
@@
Additionally, in a more general setting, one may consider an arbitrary parameterization of the Hamiltonian, for example, $H_{A}^{(\nu)}+V_{AB}^{(\nu)}+H_{B}^{(\nu)}$. In such a general framework, there may exist adiabatic paths that connect the initial ground state to an anomalous, highly entangled ground state, as demonstrated in Ref. [49]. Although we have not yet identified an explicit counterexample, we conjecture that the standard adiabatic path condition by itself is insufficient to guarantee the generalized area law.
@@
Let us denote the initial-state entanglement by the $\infty$-Rényi entanglement:
@@
Moreover, the entanglement entropy $E_{\alpha=1}(\Omega^{(1)})$ for the ground state $|\Omega^{(1)}\rangle$ is upper-bounded by
@@
In our setup, the boundary size is roughly estimated as $\tilde{g}/\Delta$ since $\tilde{g}$ characterizes the dynamical entanglement rate from the spectral SIE. Indeed, in a graph system with finite-range interactions, $\tilde{g}$ is indeed proportional to the boundary size between $A$ and $B$. From this perspective, our result qualitatively yields
@@ VII.1 Approximate-Ground-Space-Projection (AGSP)
In the proof of Theorem 4, we combine the quantum adiabatic theorem [55] with the so-called approximate ground-state projection (AGSP) approach [39, 40, 44, 43].
@@
    In the standard AGSP framework, the primary focus is to bound the Schmidt rank of the AGSP operator.
@@
For any Hamiltonian in the form of $H=H_{A}+H_{B}+V_{AB}$, there exists an AGSP operator $K_{\beta}$ with the following properties:
@@
where $\beta$ is arbitrarily chosen, and $\Delta$ has been defined as the spectral gap.
@@
Here, the AGSP $K_{\beta}$ itself may not be approximated by an operator with the Schmidt rank $D$ (see also Sec. VI). However, in applying the AGSP formalism to the ground state, the Schmidt rank approximation with respect to $K_{\beta}|\phi\rangle$ is more crucial. We use this proposition to achieve the ground state approximation in Theorem 4, as well as in Theorem 5 below, concerning the polynomial complexity of the ground state in 1D long-range interacting systems.
@@ VII.1.1 Proof of Proposition 14
In the proof, as a candidate for the AGSP operator $K_{\beta}$, we employ the following one:
@@
Next, we estimate the SE strength $\bar{{\mathcal{J}}}(K_{\beta})$ based on Definition 1 to prove the inequality (231). We use Lemma 4 and obtain the inequality
@@ VII.2 Proof of Theorem 4
We begin by considering the Schmidt-rank truncation problem for the initial ground state $|\Omega^{(\nu)}\rangle$ with $\nu=0$. In general, any restrictions to Rényi entanglement with $\alpha\geq 1$ cannot ensure the efficiency guarantee of the truncation [30, 31]. However, by utilizing the fact that $|\Omega^{(0)}\rangle$ is the gapped ground state, we can prove the following proposition, which plays a key role in the proof:
@@
$|\Omega^{(\nu)}\rangle=\sum_{s}\lambda_{s}|\phi_{A,s}\rangle\otimes|\phi_{B,s}\rangle.$ (242)
Then, we have the following general equation:
@@
where we use $\log\left(\lambda^{2}_{1}\right)=E_{\infty}(\Omega^{(\nu)})=S_{0}$. We cannot get further information on $\lambda_{s}$ with $s\geq 2$ only from the information on the $\infty$-Rényi entanglement.
To circumvent the difficulty stemming from this limitation, we recover the decay rate of $\{\lambda_{s}\}_{s}$ from the condition (243) using the AGSP formalism. To proceed along this line, we decompose $|\phi_{A,1},\phi_{B,1}\rangle$ in Eq. (242) into
@@
where $\langle\Omega^{(\nu)}|\Omega_{\bot}^{(\nu)}\rangle=0$. To recover $|\Omega^{(\nu)}\rangle$ from $|\phi_{A,1},\phi_{B,1}\rangle$, we utilize the AGSP $K_{\beta}$ constructed in Proposition 14, which yields
@@
where we use the parameter $\tilde{g}$, which upper-bounds the SE strength as in (219). Therefore, by letting $|\psi_{D}\rangle:=\lambda_{1}^{-1}|\phi_{D}\rangle$ and combining the inequalities (VII.2) and (246), we obtain
$\left\||\Omega^{(\nu)}\rangle-|\phi_{D}\rangle\right\|$ $\leq\frac{3}{\lambda_{1}}e^{-\beta\Delta^{2}}+\frac{e^{2\beta\Delta\tilde{g}}}{\lambda_{1}D^{1/2}}.$ (247)
By choosing $\beta$ such that $e^{-\beta\Delta^{2}}=e^{2\beta\Delta\tilde{g}}D^{-1/2}$, or equivalently,
@@
where we use Eq. (243) and the notation $\kappa_{\Delta}$ introduced in Eq. (241). This completes the proof. $\square$
@@
To analyze the target ground state $|\Omega^{(\nu)}\rangle$ with $\nu=1$, we next utilize the adiabatic time evolution using the parametrized Hamiltonian $H^{(\nu)}$ ($0\leq\nu\leq 1$) as follows:
@@
From Ref. [55], we obtain the approximation error as
@@
We next estimate the lower bound of $a_{\ell}$. Applying the Eckart–Young theorem [72] and employing the quantum state $|\psi_{D}\rangle$ described in Proposition 15, we obtain
@@
By combining the error bound (252) and Eq. (255) with $a_{\ell}^{2}\geq 1/2$, we derive the inequality
@@
We then construct an approximation of the target ground state $|\Omega^{(1)}\rangle$ based on the state $|\psi_{\ell,1/\epsilon}\rangle$. For this purpose, we first decompose
@@
$\sum_{s=1}^{\infty}\tilde{\lambda}_{\beta,s}\leq\sum_{s=1}^{\infty}\tilde{\lambda}_{s}\bar{{\mathcal{J}}}\left(K_{\beta}/c_{1}\right)$ $=\bar{{\mathcal{J}}}\left(K_{\beta}/c_{1}\right)e^{E_{\alpha=1/2}(\psi_{\ell,1/\varepsilon})/2}\leq 3\ell^{1/2}e^{2\beta\Delta\tilde{g}+\tilde{g}/\varepsilon},$ (267)
@@
$\left\|c_{1}^{-1}K_{\beta}|\psi_{\ell,1/\epsilon}\rangle-|\psi_{\beta,D}\rangle\right\|\leq\frac{3\ell^{1/2}e^{(2\beta\Delta+1/\varepsilon)\tilde{g}}}{D^{1/2}},$ (268)
where we use an inequality similar to (33).
@@
By choosing the parameter $\beta$ such that $e^{-\beta\Delta^{2}}=e^{2\beta\Delta\tilde{g}}D^{-1/2}$, or equivalently,
@@
Finally, we upper-bound the coefficient $12\ell^{1/2}e^{\tilde{g}/\varepsilon}$ in Eq. (271). As in Eqs. (258) and (261), $\ell$ and $\varepsilon$ have the following upper and lower bounds, respectively:
@@
$\left\||\psi\rangle-|\psi_{D}\rangle\right\|\leq\frac{C}{D^{\kappa}}\quad\forall D\in\mathbb{N},$ (274)
where $C$ and $\kappa$ are positive constants. Then, the entanglement entropy is upper-bounded by
@@
$c_{\kappa,1}:=\frac{2-2^{-2\kappa}}{\kappa(1-2^{-2\kappa})},\quad c_{\kappa,2}:=\frac{(6+2\kappa)\log(2)}{(1-2^{-2\kappa})^{2}}.$ (276)
By applying this lemma with the first main inequality (223), i.e., $\kappa\to\kappa_{\Delta}$ and $C\to C_{\tilde{g},\Delta,S_{0}}$, we immediately prove the second main inequality (226). This completes the proof of Theorem 4. $\square$
Proof of Lemma 16. We consider a set of $\{D_{p}\}_{p=1}^{\infty}$ such that $D_{p}=2^{p+p_{0}}$ and define $\delta_{p}$ as
@@
where the value of $p_{0}$ will be chosen appropriately to satisfy a condition required for the proof. Temporarily, we leave $p_{0}$ free and proceed. Then, by using Ref. [44, Supplemental Proposition 3], we obtain
@@
$\leq p_{0}\log(2)+\frac{2^{-2p_{0}\kappa}C^{2}}{1-2^{-2\kappa}}\log\left(\frac{6\cdot 2^{p_{0}(2\kappa+1)}}{C^{2}}\right)+\frac{C^{2}2^{-2\kappa(p_{0}+1)}(2\kappa+1)\log(2)}{(1-2^{-2\kappa})^{2}},$ (278)
@@
VIII Simulating 1D long-range interactions with polynomial complexity